OpenStudy (anonymous):

27. Write the equation of a line that passes through the points ( -2, 5) and (-2, 11) @MathLegend

OpenStudy (mathlegend):

First, we need a slope. y2-y1/x2-x1

OpenStudy (mathlegend):

Can you @kendoll substitute in the proper values?

OpenStudy (anonymous):

11-5/-2-2 ?

OpenStudy (mathlegend):

11-5/-2+2

OpenStudy (anonymous):

6/0?

OpenStudy (mathlegend):

Yes.

OpenStudy (anonymous):

thts it ?

OpenStudy (mathlegend):

No.

OpenStudy (anonymous):

so wat we do now

OpenStudy (mathlegend):

Now, we need to use... point slope form... y-y1=m(x-x1) Can you substitute in those values?

OpenStudy (anonymous):

no?

OpenStudy (mathlegend):

y-5=0(x+2)

OpenStudy (mathlegend):

Can you solve for y?

OpenStudy (anonymous):

y-5=0(x+2) +5 +5

OpenStudy (mathlegend):

ok so what does y equal?

OpenStudy (anonymous):

y=5(x+2)?

OpenStudy (mathlegend):

@kendoll y-5=0(x+2) Remember anything multiplied by 0 is 0... y-5 = 0 +5 +5 ---------- y = 5

OpenStudy (mathlegend):

y= 5 is your final answer. It can not be y=mx+b since a slope of zero is just a horizontal line.

OpenStudy (mathlegend):

If this were graphed it would look like...

OpenStudy (anonymous):

wat about the x+2 it dissapear

OpenStudy (mathlegend):